OpenStudy (anonymous):

integrate : dx/square root 28-12x+x^2

OpenStudy (turingtest):

\[\int{dx\over28-12x+x^2}\]?

OpenStudy (lgbasallote):

i think there should be a swrt

OpenStudy (anonymous):

\[dx \div \sqrt{28-12x+x^2}\]

OpenStudy (anonymous):

use partial fraction method.....should work

OpenStudy (turingtest):

oops\[\int{dx\over\sqrt{28-12x+x^2}}\]

OpenStudy (lgbasallote):

hmm looks like trig sub though

OpenStudy (turingtest):

definitely trig sub

OpenStudy (anonymous):

yes thats right Turing test its integration by parts

OpenStudy (lgbasallote):

arcsin?

OpenStudy (lgbasallote):

oh by parts

OpenStudy (anonymous):

What if we write the bottom as (x-6)^2 - 8 and then u = x - 6?

OpenStudy (anonymous):

wow again? lol

OpenStudy (anonymous):

i've answered 2 integration by parts today already >_>

OpenStudy (anonymous):

So we are left with \[\int \frac{dx}{\sqrt{(x-6)^2 - 8}} = \int \frac{du}{\sqrt{u^2 - 8}}\]

OpenStudy (turingtest):

..and how is that not a trig sub?

OpenStudy (anonymous):

Afterwards, yes, I agree. So disregard.

OpenStudy (turingtest):

I was trying to imagine a PF way to do it... got me all interested haha

OpenStudy (turingtest):

@roselynECE do you know what substitution to make?

OpenStudy (anonymous):

this is not substitution its integration by parts

OpenStudy (turingtest):

are you sure about that?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

integrals with fractional exponents don't usually involve integration by parts unless it's got a logx with it or something, so good luck to whoever figures out how to do this by parts!

OpenStudy (anonymous):

i see inverse trig

OpenStudy (turingtest):

me too

OpenStudy (turingtest):

...which can be shown with a trig sub...

OpenStudy (anonymous):

Im sorry for that I was just overlook

OpenStudy (turingtest):

no apologies necessary :)

OpenStudy (anonymous):

Haha @TuringTest sorry for that. At first glance, I thought I could expand it out into PF. But, yeah, hopes were gone.

OpenStudy (anonymous):

how to solve this using arcsin?

OpenStudy (anonymous):

What I did from above is to say that u = sqrt(8)*sec(v) and du = sqrt(8)*tan(v)sec(v)dv but I suppose there are easier ways.

OpenStudy (lgbasallote):

if you use completing the square...youll get \(\LARGE \int \frac{dx}{\sqrt{(6-x^2)^2 - 8}}\)

OpenStudy (lgbasallote):

that's (6-x)^2 in the denom

OpenStudy (lgbasallote):

draw a triangle...let the hypotenuse = 6- x.base leg = \(\sqrt 8\) and the other leg = \(\sqrt{(6-x)^2 - 8}\) so let \(\sqrt 8 \sec \theta = 6-x\) \(\sqrt 8 \tan \theta = \sqrt{(6-x)^2 - 8}\) \(\sqrt 8 \sec \theta \tan \theta d\theta = -dx\) substitute these values to your integral

OpenStudy (lgbasallote):

can you tell me what you have from substituting those into your integral @rocelynECE

OpenStudy (lgbasallote):

@roselynECE sorry misspelled last time